OpenStudy (anonymous):

What is the solution to p/.3 – 1 = 5?

OpenStudy (anonymous):

p = 1.8

OpenStudy (anonymous):

Are you positive?

OpenStudy (anonymous):

is it (p/0.3)-1 = 5? or p/(0.3-1) = 5 if its the first one, i am sure. you just multiply everything by 0.3 then solve.

OpenStudy (anonymous):

there isn't any parentheses

OpenStudy (anonymous):

well i just put the parentheses to split up the numbers, but if the are no parentheses, then it looks like my first answer is right.

OpenStudy (anonymous):

hmmm then why did I keep getting 1.2? I know 1.2 is wrong but I can't get anything else...

OpenStudy (anonymous):

you must have subtracted o.3 from each side, but you had to add. because when you multiply everything by 0.3, you end up with p-0.3=1.5. then you add o.3 to each side, not subtract. so the answer would be 1.8.

OpenStudy (anonymous):

no I didn't. I subtracted 1 from both sides, them multiplied both sides by .3.... Idk... I will just go over it with my teacher tomorrow...

OpenStudy (anonymous):

you are supposed to add one to each side, not subtract.

OpenStudy (anonymous):

well that explains everything :) LOL! thanks! :)

OpenStudy (anonymous):

no problem :)